Circular No. 8644

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V477 SCUTI S. Mazuk, D. K. Lynch, R. J. Rudy, and C. C. Venturini, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; R. B. Perry, Langley Research Center, NASA; and B. Walp, Lick Observatory, report 0.47- to 2.5-μm spectroscopy of V477 Sct (cf. IAUC 8617) with the Lick 3.0-m telescope (+ VNIRIS) at Nov. 15.094 UT. The nova showed doubled line profiles with FWHM of 2700 km/s. The Lyβ-fluoresced O I lines were strong, as were those of H I and He I. He II was present but weak. Coronal lines of [S VIII] and [S IX] were evident, but not those of [Ca VIII] or [Si VI]. As derived from the O I lines, the reddening is E(B − V) = 1.2. A strong continuum increasing toward longer wavelengths appears to be thermal emission from dust.

COMETS C/2005 T9–T11 AND C/2005 U2–U6 (SOHO) Further to IAUC 8640 (where on line 4 the designation should read

C/2005 S13), additional near-sun comets have been found on SOHO website images — all the following being Kreutz sungrazers except for C/2005 T9 (Meyer group). C/2005 T9 was tiny, stellar, and too faint for photometry. C/2005 T10 was very faint and diffuse. C/2005 T11 reached mag 4.9 at 10.9R_{⊙} on Oct. 15.821 UT, as seen in C3 images; in C2 images, it showed a bright, long, thin tail that reached a length of ∼ 38′ at 3.5R_{⊙} on Oct. 16.226. C/2005 U2 and C/2005 U3 were small, very faint, and stellar in appearance in C3 images, while they were very faint and diffuse in C2 images. C/2005 U4 was diffuse and very faint. C/2005 U5 was small, very faint, and stellar in appearance. C/2005 U6 was tiny, faint, and stellar in appearance in C3 images, and it appeared as a small, faint blob in C2 images.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 T9 Oct. 8.226 12^{h}52.^{m}3 − 4^{o}00′ C2  HS 2005-X52
C/2005 T10 14.726 13 11.8 − 9 13 C2 HS 2005-X52
C/2005 T11  15.029 12 57.3 −11 39 C3/2 TH 2005-X52
C/2005 U2  18.113 13 19.3 −11 00 C3/2 HS 2005-X53
C/2005 U3  18.513 13 14.5 −11 32 C3/2 HS 2005-X53
C/2005 U4  19.454 13 30.2 −11 04 C2  HS 2005-X53
C/2005 U5  21.448 13 30.6 −13 36 C3  TS 2005-X53
C/2005 U6  23.447 13 39.1 −13 13 C3/2 BZ 2005-X53
2005 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green